# Explain What You Mean: Intent Augmented Knowledge Graph Recommender Built With An LLM

### v2

@@ 4.4. Intent Translation GNN (Q2)
(8) $\mathbf{r}^{u,i}=\text{softmax}(\mathbf{P}^{u,i})^{T}\mathbf{D}^{u,i}$
4.5. Intent Aware Scoring (Q3)
@@
(9) $z_{u,i}=\frac{\mathbf{e}_{u}\cdot\mathbf{e}_{i}}{\parallel\mathbf{e}_{u}\parallel\cdot\parallel\mathbf{e}_{i}\parallel}$
We denote a single intent extracted from $q$ as $n^{(q)}$, and the collection of all intents for $q$ as $\Omega_{q}$. Similarly, a single intent extracted from $u$ and their collections is denoted as $n^{(u)}$ and $\Omega_{u}$, and a single intent extracted from $i$ and their collections is denoted as $n^{(i)}$, and $\Omega_{i}$. Additionally, $\mathbf{e}_{x}$ is used to represent the embedding of some entity $x$.
@@
(10) $y_{u,i}=\mathop{max}\limits_{n^{(u)}\in\Omega_{u},n^{(i)}\in\Omega_{i}}\frac{\mathbf{e}_{n^{(u)}}\cdot\mathbf{e}_{n^{(i)}}}{||\mathbf{e}_{n^{(u)}}\parallel\cdot\parallel\mathbf{e}_{n^{(i)}}\parallel}\cdot 0.5^{\mathbb{I}(\Omega_{u}\cap\Omega_{i}=\varnothing)}$
Finally, the score for the triplet $(u,i)$ is the hybrid mixture of the embedding score and intent matching score:
(11) $score(u,i)=y_{u,i}+\lambda z_{u,i}$
@@
where $\lambda$ is an empirical mixer coefficient.
4.6. Inference Procedure And Complexity Analysis
5. Experiments
We use Llama-3.1-8B model for LLM inference, and all-mpnet-base-v2 from sentence transformer for encoding textual features. We retrieve the top 100 prebuilt intent candidates in the kNN retrieval step of RAG. We apply an 8:1:1 ratio when sampling the positive/negative edges for train/dev/test sets in the graph. The coefficient $\lambda$ in Equation 11 is set to 0.1 based on grid search result on the validation set.
[…]
Datasets Search Beauty Books Steam Yelp2022
#Users 36 0333603336\,03336 033 40 2264022640\,22640 226 251 394251394251\,394251 394 281 428281428281\,428281 428 1 987 89819878981\,987\,8981 987 898
#Items 872 678872678872\,678872 678 54 5425454254\,54254 542 25 6062560625\,60625 606 13 0441304413\,04413 044 150 347150347150\,347150 347
#Inter 3.5M 0.35M 3.2M 3.5M 6.9M
Density 0.011% 0.02% 0.05% 0.095% 0.002%
#Intents 495 285495285495\,285495 285 39 3053930539\,30539 305 45 9324593245\,93245 932 53 9405394053\,94053 940 65 0406504065\,04065 040
#IntEdges 6.9M 209K 390K 231K 409K
AvgIntDeg 13.9 5.3 8.5 4.3 6.3
5.1. Baselines and Datasets
Enterprise Search. In the enterprise search setting, a set of developers query the search engine to search for datasets published by other developers within the enterprise. The search engine first uses BM25 (Whissell and Clarke, 2011) to retrieve a list of item candidates, then the IKGR is applied to rerank the search results. Evaluations are done over user’s feedback on how high the item of interest could be ranked among the final reranked list. In this setting, the items are enterprise datasets, which consist of text features such as dataset name, description, column names, ID labels, etc. The users are all registered within the enterprise and their developer profiles could be collected through separate channels to offer a hint about their dataset consumption preference. The IKGR is trained over historical user-dataset consumption interactions collected from separate channels. This dataset is labeled as Search in dataset description table and result tables.
Books, Beauty. These datasets are obtained from Amazon review [^1] http://jmcauley.ucsd.edu/data/amazon/ in (McAuley et al., 2015), which contains a variety of categories. We utilize the Books and Beauty categories. We leverage the features of title, sales type, sales rank, categories, price, and brand.
Steam [^2] https://cseweb.ucsd.edu/~jmcauley/datasets.html#steam_data. This is a dataset collected from Steam (Kang and McAuley, 2018), a large online video game distribution platform. We leverage the item features of app name, genres, publisher, sentiment, specs, tags.
To evaluate the ranking list of all the models, we employ Hit Ratio (HR), Normalized Discounted Cumulative Gain (NDCG), and Mean Reciprocal Rank (MRR). Considering we only have one ground truth item for each user, HR@$k$ is equivalent to Recall@$k$ and proportional to Precision@$k$; MRR is equivalent to Mean Average Precision (MAP). We report HR and NDCG with $k={1,5,10}$. For all these metrics, the higher the value, the better the performance.
To verify the effectiveness of our method, we compare it with the following representative baselines:
POP: It is the simplest baseline that ranks items according to their popularity judged by the number of interactions.
BPR-MF (Rendle et al., 2012): It optimizes the matrix factorization with implicit feedback using a pairwise ranking loss.
FM (Rendle, 2010): The factorization machine models all interactions between variables using factorized parameters.
ConvNCF (He et al., 2018): It employs a convolutional neural network to learn high-order correlations among embedding dimensions neural collaborative filtering.
FPMC (Rendle et al., 2010): It captures users’ general taste as well as their sequential behaviors by combing MF with first-order Markov chains.
RippleNet (Wang et al., 2018): It uses a single layer of embedding translation by applying transH to user-item interaction.
KGCN (Wang et al., 2019): The KG based graph convolutional networks captures inter-item relatedness effectively by mining their associated attributes on the KG.
KTUP (Cao et al., 2019): It uses a single layer of embedding translation by applying TransH to user-item interaction.
Table 2. Performance comparison of different methods. Bold scores are the best in each row, while underlined scores are the second best.
Datasets Metric POP BPR-MF ConvNCF FPMC RippleNet KGCN KTUP IKGR
Search HR@1 0.0001 0.0041 0.0075 0.0048 0.0024 0.0051 0.0078 0.0086
HR@5 0.0023 0.0100 0.0149 0.0125 0.0118 0.0161 0.0166 0.0202
HR@10 0.0076 0.0174 0.0262 0.0225 0.0218 0.0232 0.0262 0.0267
NDCG@5 0.0029 0.0068 0.0090 0.0074 0.0068 0.0101 0.0137 0.0151
NDCG@10 0.0046 0.0094 0.0112 0.0107 0.0087 0.0123 0.0142 0.0172
MRR 0.0046 0.0108 0.0070 0.0097 0.0091 0.0100 0.0130 0.0153
Beauty HR@1 0.0405 0.0552 0.1005 0.1009 0.0532 0.0984 0.1783 0.1369
HR@5 0.0377 0.1844 0.2754 0.2379 0.1972 0.3120 0.3388 0.3316
HR@10 0.1074 0.3540 0.4387 0.3957 0.3695 0.4204 0.4610 0.4846
NDCG@5 0.0763 0.1245 0.1758 0.1335 0.1307 0.1729 0.2583 0.2806
NDCG@10 0.0443 0.1731 0.1961 0.2123 0.1713 0.2345 0.2814 0.2939
MRR 0.0909 0.1792 0.1736 0.1752 0.1382 0.2254 0.2581 0.2641
Books HR@1 0.0015 0.0324 0.1124 0.0620 0.0480 0.0853 0.1125 0.1251
HR@5 0.0524 0.1663 0.2248 0.1859 0.1691 0.2465 0.2652 0.3197
HR@10 0.1073 0.2649 0.3181 0.3217 0.3553 0.3582 0.3614 0.4248
NDCG@5 0.0314 0.0983 0.1371 0.1268 0.1031 0.1579 0.1921 0.2097
NDCG@10 0.0532 0.1603 0.1633 0.1703 0.1607 0.2016 0.2415 0.2814
MRR 0.0788 0.1317 0.1567 0.1532 0.1407 0.1832 0.2193 0.2672
Steam HR@1 0.0091 0.0547 0.0860 0.0526 0.0304 0.0641 0.1060 0.1095
HR@5 0.0370 0.1196 0.1976 0.1698 0.1429 0.1966 0.2445 0.2759
HR@10 0.0901 0.2379 0.3374 0.2717 0.2735 0.3218 0.3126 0.3574
NDCG@5 0.0291 0.1035 0.1121 0.1094 0.0915 0.1496 0.1614 0.1735
NDCG@10 0.0459 0.1244 0.1471 0.1384 0.1161 0.1905 0.1809 0.2212
MRR 0.0507 0.1265 0.1382 0.1272 0.1294 0.1528 0.1957 0.2168
Yelp2022 HR@1 0.0171 0.0636 0.1177 0.0524 0.0495 0.0917 0.1005 0.0989
HR@5 0.0366 0.1555 0.1987 0.1911 0.1513 0.2405 0.2395 0.2869
HR@10 0.1018 0.2462 0.3576 0.3028 0.2996 0.3395 0.3317 0.3966
NDCG@5 0.0233 0.0823 0.1395 0.1131 0.1201 0.1397 0.2067 0.2078
NDCG@10 0.0334 0.1364 0.1434 0.1527 0.1225 0.1861 0.2145 0.2277
MRR 0.0593 0.1327 0.1542 0.1200 0.1356 0.1578 0.1906 0.2100
From the approaches perspective, IKGR shows new state-of-the-art performances especially on the Search dataset with the largest knowledge gap. On the other hand, across these datasets, the knowledge graph embedding translation approaches (IKGR, TKUP) or GNN based approaches (KGCN, RippleNet) performs better than factorization machines or collaborative filtering based approaches (BPR-MF, FPMC, ConvNCF), demonstrating the gain from structural embedding learning.
5.2. Ablation Studies
Table 3. Ablation Study
Metrics HR@1 HR@5 HR@10 NDCG@5 NDCG@10 MRR
IKGR 0.0086 0.0202 0.0267 0.0151 0.0172 0.0153
Exact Int. Only 0.0077 0.0185 0.0254 0.0147 0.0163 0.0134
No Aug. 0.0073 0.0175 0.0240 0.0131 0.0150 0.0125
Inter. Aug. 0.0070 0.0166 0.0220 0.0126 0.0141 0.0117
No Int. scoring 0.0082 0.0193 0.0256 0.0140 0.0164 0.0139
No Int. Trans. 0.0078 0.0184 0.048 0.0132 0.0152 0.0130
To investigate the influence of each component over the enterprise search data, we conduct a list of experiments that drop each component in the proposed pipeline and compare with the original pipeline on the Search dataset. Results are presented in Table 3.
In Table 3, IKGR means the original approach without component dropping, Exact Int. Only means dropping the second round of related intent retrieval using RAG, and only exact intent edges are presented in the graph, without the related intent edges. No Aug. means training the architecture of IKGR, which is a naive knowledge graph convolution followed by an embedding translation layer, without adding the intent node in the graph. Inter. Aug. means to employ LLM only as a user-item implicit feedback augmentor without adding the intent nodes (Wei et al., 2024). In this setting the LLM is asked to select items that the user might be likely or unlikely to interact with. No Int. scoring means using a simple cosine similarity based score and we do not apply intent-aware scoring function as in Equation 11. No Int. Trans. means intent translation is not leveraged to compute the translation as in Equation 8.
As observed in the results shown in Table 3, the relative contributions of each components of IKGR can be estimated. The most significant observation is that the intent edges augmentation steps as a whole helps to densify the graph almost twice, as reflected by the #IntEdges and #Inter in Table 1, and the MRR performance correspondingly improved 22% (0.0125 to 0.0153) shown in Table 3. This confirms that adding intent relations to the graph improves the IKGR performance while densifying the graph with meaningful knowledge connectivities.
6. Conclusions

### v3

@@ B.2 Intent Prior Injected Translation layers
$\mathbf{r}^{u,i}=\text{softmax}(\mathbf{P}^{u,i})^{T}\mathbf{D}^{u,i}$ (8)
B.3 Intent Aware Scoring
@@
$z_{u,i}=\frac{\mathbf{e}_{u}\cdot\mathbf{e}_{i}}{\parallel\mathbf{e}_{u}\parallel\cdot\parallel\mathbf{e}_{i}\parallel}$ (9)
We denote a single intent extracted from item $i$ as $n^{(i)}$, and the collection of all intents for $i$ as $\Omega_{i}$. Similarly, a single intent extracted from $u$ and their collections is denoted as $n^{(u)}$ and $\Omega_{u}$. Additionally, $\mathbf{e}_{n}$ is used to represent the embedding of some entity $n$.
@@
$y_{u,i}=\mathop{max}\limits_{n^{(u)}\in\Omega_{u},n^{(i)}\in\Omega_{i}}\frac{\mathbf{e}_{n^{(u)}}\cdot\mathbf{e}_{n^{(i)}}}{||\mathbf{e}_{n^{(u)}}\parallel\cdot\parallel\mathbf{e}_{n^{(i)}}\parallel}\cdot 0.5^{\mathbb{I}(\Omega_{u}\cap\Omega_{i}=\varnothing)}$ (10)
Equation 10 has two components, the similarity component and the non-overlap punishment component. In the event that $u$ and $i$ share a single intent (no need to have completely equal intent set), the non-overlap punishment component will be disabled (=1), otherwise if no intents are shared, it will punish by 0.5. The choice of 0.5 is decided based on empirical comparison along the development of the approach, which shows priority over too severe punishment values (e.g. 0.1). Finally, the score for the triplet $(u,i)$ is the hybrid mixture of the embedding score and intent matching score:
$score(u,i)=y_{u,i}+\lambda z_{u,i}$ (11)
@@
where $\lambda$ is an empirical mixer coefficient. In the experiments, the coefficient $\lambda$ is set to 0.1 based on grid search result on the validation set.
In the ablation study section Section 5.3, Opt2. Vanilla GNN means translation layers, simply use a two-layer GNN trained with standard ranking loss. Opt3. Vanilla Trans. means simply use a translation layer without GNN and the scoring function of Equation 11. Opt4. Vanilla Scoring means keep both GNN and translation layer, but removing intent-aware scoring function of Equation 11.
Appendix C Related Works Continued
Recommender Approaches Addressing Data Sparsity. Recent works tackle the data sparsity issue through enhanced model architectures, graph-based techniques, and knowledge-aware methods. The Item History Model (IHM) improves cold-start item recommendations by directly injecting user-interaction data into the item tower and employing an inductive structure for real-time inference [6]. Cold Brew addresses sparsity in graph neural networks by distilling node representations, mitigating the impact of missing or noisy neighbors [35]. RippleNet further alleviates sparsity by propagating user preferences through knowledge graph relations, enriching item representations beyond collaborative signals [7]. These approaches demonstrate that leveraging historical interactions, distilling structural knowledge, and integrating external information effectively mitigate data sparsity in recommendation systems.
Directly Leveraging LLM as the Recommender. A-LLMRec enhances LLM-based recommendation by leveraging embeddings from state-of-the-art collaborative filtering models, excelling in both cold and warm scenarios while maintaining efficiency and model-agnostic integration [36]. CoRAL introduces collaborative retrieval-augmented prompting, addressing LLMs’ reliance on semantic information by incorporating user-item interactions through reinforcement learning-based retrieval policies, significantly improving long-tail recommendation [37]. TALLRec proposes a tuning framework to align LLMs with recommendation-specific tasks, demonstrating strong generalization and efficiency even with limited training data [38]. These approaches collectively highlight the potential of LLMs as standalone recommenders by addressing cold-start challenges, enhancing collaborative reasoning, and improving alignment with recommendation-specific objectives.
Personalized Re-ranking Recommender System. Re-ranking in recommender systems aims to refine an initially ranked list to better capture user preferences and item relationships. Traditional ranking methods optimize global performance but often overlook the mutual influence between items and user-specific intent [4]. Recent approaches address these limitations by integrating semantic tag information [39], multi-stage ranking optimization [40], and item relationships [5]. Transformer-based models effectively model global item interactions, while graph-based methods leverage item relationships for improved ranking [4, 5]. Furthermore, self-supervised learning enhances sequential recommendation by mitigating data sparsity issues [41]. These methods collectively demonstrate that incorporating user personalization, item dependencies, and efficient ranking strategies significantly enhances re-ranking effectiveness.